# Multi-GraspLLM: A Multimodal LLM for Multi-Hand Semantic Guided Grasp Generation

### v1

@@ 4.2 Data Format
Figure 5: The template of training data used to train the Multi-GraspLLM.
As shown in Figure 5, we introduce three special tokens in the Vicuna [4] backbone fine-tuning: (1) a Robotics hand special token for hand identification, (2) a Scale special token for point cloud size normalization, and (3) a Grasp special bin token mapping to 512 discretization bins.These tokens work together to ensure accurate information processing during both tokenization and de-tokenization.
With the help of these specialized tokens, we combined and refined the basic conversations from Multi-GraspSet and the caption task. The final dataset consists of four types of conversations, as illustrated in Figure 5: (1) Single-Round Grasp Generation; (2) Multi-Round Mix; (3) Multi-Round Grasp Generation. The first type involves single-round interactions, where the model predicts grasping angles based on simple queries like ”How can the {robotic hand} grasp the object?”. The latter two types feature multi-round interactions. ”Multi-Round Mix” starts with object class description and follows with grasp prediction. The ”Multi-Round Grasp Generation” supports more diverse dialogues involving different robotic hands and requiring sophisticated grasp predictions, going beyond the basic approach. We purposely incorporated captioning tasks to enhance the model’s 3D geometric understanding and reduce its reliance on text-only inputs. Our ablation studies in Sec. 5.4 validate that this strategy significantly improves performance, achieving a balanced interpretation of both spatial geometry and grasping actions.
@@ 4.3 Training Strategy
Stage1: Multimodal Alignment. In multimodal alignment, we aim to have the adaptor map point cloud modality data and language modality data into a space that is as close as possible. We freeze the LLM backbone and point cloud encoder while fine-tuning only the modality adaptor. The fine-tuning process incorporates single-round conversations ”Single-Round Grasp Generation” described in Sec. 4.2.
@@ 5 Experiments
method Allegro Shadow Panda Avg.
@@
GraspTTA [11] 1.31 1.12 0.26 0.075 1.20 1.18 0.31 0.062 / / / / 1.27 1.15 0.29 0.068
SceneDiffuser [8] 0.74 1.04 0.31 0.086 1.02 0.86 0.37 0.081 / / / / 0.88 0.96 0.35 0.084
6Dof-GraspNet [24] / / / / / / / / 0.59 0.41 0.40 / / / / /
Contact-GraspNet [34] / / / / / / / / 0.62 0.43 0.46 / / / / /
Multi-GraspLLM-mix 0.43 0.97 0.29 0.092 0.48 0.80 0.42 0.089 0.29 0.37 0.45 / 0.40 0.71 0.38 0.091
Multi-GraspLLM-split 0.39 0.99 0.32 0.094 0.45 0.72 0.40 0.083 0.26 0.42 0.48 / 0.37 0.70 0.40 0.089
Table 2: Comparison of SOTA and Multi-GraspLLM. ’-mix’ denotes models trained on mixed datasets, while ’-split’ indicates models trained separately for each robotic hand.
5.1 Dataset and Metric
@@
Multi-GraspLLM is evaluated in two key aspects: grasp intention and physical stability. To evaluate grasp intention accuracy, we employ Chamfer distance (CD) (cm), which quantifies the spatial alignment between the predicted and ground truth robotics hand point clouds. This metric measures how well our inferred grasps match the intended grasp locations on objects as specified in the descriptions, where lower values indicate better alignment. To evaluate grasp quality, we employ three metrics: the maximum penetration distance (Pen-d) (cm) between the gripper and target object measures physical feasibility, the grasp success rate (Suc.) in Isaac Sim and the Q1 metric which is intuitively the norm of the smallest wrench that can destabilize the grasp [42]. More details are in our supplementary material.
@@ 5.2 Implementation Details
For constructing Multi-GraspSet, we employ DexGraspNet [42] and Contact-GraspNet [34] as our grasp generation methods. For DexGraspNet, we generate 128 grasp poses for the Allegro Hand [45] and 160 for the Shadow Hand [31] per object, using Adam optimizer with 6000 epochs. For Contact-GraspNet [34], we quickly sample 256 points from each point cloud and generate 200 grasps.
During training, we perform multimodal alignment for 3 epochs with a learning rate of 2e-3, followed by instruction tuning for another 3 epochs with a learning rate of 2e-5. For each object, we uniformly sample 8,192 points from its mesh to generate the point cloud, which is then tokenized into 512 tokens. The grasp angles are discretized into 384 bins, with their upper and lower bounds for different robotic hands determined by their entire dataset.
@@ 5.3 Comparison with State-of-the-Art Methods
We compared our Multi-GraspLLM method with existing open-source SOTA approaches. Since there is currently no open-source method for multi-dexterous hand and gripper grasping generation, we conducted separate comparisons on corresponding end-effectors: SceneDiffuser [8] and GraspTTA [11] were evaluated on dexterous hands, both trained on our dataset, while Contact-GraspNet [34] and 6dof-GraspNet [24] were assessed on parallel grippers. The main results are shown in Table 2, ”GraspLLM-split” means our Multi-GraspLLM model trained on the dataset separately for each manipulator, while ”-mix” refers to mixed training. We can see that mixed training performs slightly worse than separate training, but the drop in performance is minor. Whether trained separately or mixed, both methods outperform all existing baselines, especially in intention accuracy, with better physical stability and less penetration. Unlike DexGraspNet [42], which adds specific loss functions for physical stability, and the two-stage method ”Grasp as You Say” [44], our approach is end-to-end without any extra loss functions. Although the physical stability of our method does not exceed the limits of the training dataset, it still outperforms all other compared methods.
As shown in Figure 6, Multi-GraspLLM can generate different robotic hand grasping poses based on instructions with varying levels of contact information. Our approach supports grasp generation at different instruction levels. For the low-level instruction, model generates grasps without any contact information, while mid-level focuses on the object part information. The high-level instruction utilizes finger contact information, enabling our model to simply control individual fingers.
@@ 5.4 Ablation Study
To verify the effectiveness of each component of our method, we conducted ablation studies, examining the impact of each special token and the presence of two-stage training. Additionally, we explored the effects of grasp bin length and explored how the structure and size of training data affect the model’s performance.
@@
Multi-GraspLLM 0.40 0.71 0.38 0.091
w/o hand token 0.43 0.78 0.38 0.088
w/o scale 0.49 1.01 0.32 0.081
w/o grasp 0.45 1.09 0.32 0.075
w/o 2-stage 0.47 1.11 0.34 0.071
@@
We first performed ablation on each special token. As shown in Table 3, the two-stage training and the existence of grasp bin token significantly impact performance. A single-stage model struggles to respond effectively to newly introduced grasp bin tokens it has not encountered before. Due to the incomplete encoding vocabulary for numbers in Vicuna’s [4] tokenizer, larger numbers without grasp bin token such as ”400” may be decoded as separate tokens [”40”,”0”], which further hinders the model’s grasp bin comprehension. Additionally, since PointBERT [52] normalizes point clouds, providing a scale token helps the model better understand point cloud shapes. Lastly, the robotic hand special token has minimal impact in our case, as we only use three types of robotic hands, which the model can memorize by name. However, with a larger variety of hands, the robotic hand special token would play a more significant role.
@@
Next, we examined the effect of the number of grasp bins. The choice of bin number presents a trade-off: smaller bin numbers (resulting in larger bin widths) provide higher resolution, but make prediction more challenging. Larger bin numbers simplify prediction but reduce the precision of multi-hand reconstruction. We therefore selected a moderate bin width to balance these competing factors.
@@
We also analyzed the effect of the size of the training data in modality alignment and structure of data in instruction tuning, as shown in Figure 7 and Table 5. We gradually reduced the amount of modality alignment data to test our alignment method. As shown in Figure 7, using more data improves the performance of our model in intention understanding (lower CD) and grasp quality (higher Q1 and Suc.).
We then performed an ablation study on the composition of the instruction tuning data. As described in Section 4.2, our instruction tuning data consists of three types. ”Single-Grasp” represents the ”Single-Round Grasp Generation”, and ”Mix” is the abbreviation of ”Multi-Round Mix”. ”Multi-Grasp” represents multi-round conversations for generating grasps, corresponding to ”Multi-Round Grasp Generation” in Section 4.2. As shown in Table 5, more complex and multi-round conversations used during fine-tuning lead to better performance. Due to the introduction of caption data, the model can be prevented from excessively focusing all attention on natural language while ignoring the geometric features of point clouds, which leads to improved grasp prediction intention.
@@
0.43 0.73 0.35 0.086
0.42 0.76 0.36 0.089
0.40 0.71 0.38 0.091
Table 5: Ablation on training data for instruction tuning.
6 Conclusion
In this work, We present the first multi-hand grasp dataset with semantic guidance, along with a multimodel LLM for multi-hand grasp generation named Multi-GraspLLM. It can generate grasp poses for different robotic hands based on linguistic grasp descriptions. Our approach fills a gap in the cross-hand semantic grasp generation field. In the future, we plan to collect larger data using more types of robotic hands and objects, which can further enhance the ability of Multi-GraspLLM to generalize to real-world scenarios, ultimately driving advancements in multi-robot collaboration and human-robot interaction.

### v3

@@ 4.2 Data Format
Figure 5: The template of training data used to train Multi-GraspLLM.
As shown in Figure 5, we introduce three special tokens in the Vicuna [5] backbone fine-tuning: 1) Robotics hand special token for hand identification, 2) Scale special token for point cloud size normalization, and 3) Grasp special bin token mapping to 512 discretization bins. These tokens work together to ensure accurate information processing during both tokenization and de-tokenization.
With these specialized tokens, we combined and refined the basic conversations from Multi-GraspSet and the caption task. The final dataset consists of four types of conversation, as shown in Figure 5: 1) Single-Round Grasp Generation, 2) Multi-Round Mix, and 3) Multi-Round Grasp Generation. The first type involves single-round interactions, where the model predicts grasping angles based on simple queries like “How can the {robotic hand} grasp the object?” The latter two types feature multi-round interactions. “Multi-Round Mix” starts with object class description and follows with grasp prediction, and “Multi-Round Grasp Generation” supports more diverse dialogues involving different hands and requiring sophisticated grasp predictions, going beyond the basic approach. We purposely incorporated captioning tasks to enhance the model’s 3D geometric understanding and reduce its reliance on text-only inputs. Our ablation studies in Sec. 5.4 validate that this strategy significantly improves performance, achieving a balanced interpretation of spatial geometry and grasping actions.
@@ 4.3 Training Strategy
Stage 1: Multimodal Alignment. In multimodal alignment, we aim to align the point cloud embedding space as closely as possible with the language feature space through the adaptor. We freeze the LLM backbone and point cloud encoder while fine-tuning only the modality adaptor. The fine-tuning process incorporates single-round conversations “Single-Round Grasp Generation” described in Sec. 4.2.
@@ 5 Experiments
Model Allegro Shadow Barrett Jaco Panda Avg
@@
GroundTruth 0.65 1.08 0.29 1.07 0.79 0.38 1.09 0.43 0.31 1.38 0.59 0.37 0.65 0.41 0.44 0.97 0.66 0.36
DexGraspNet [49] 0.62 1.27 0.31 0.94 0.83 0.35 1.30 0.63 0.21 1.20 0.45 0.22 / / / 1.02 0.80 0.28
SceneDiffuser [9] 0.74 1.04 0.31 1.02 0.86 0.34 1.21 0.72 0.24 1.58 0.71 0.29 / / / 1.14 0.76 0.30
GraspNet [41] / / / / / / / / / / / / 0.62 0.43 0.46 / / /
DexGYS [52] 0.5 1.18 0.27 0.54 0.83 0.39 0.78 0.29 0.29 0.67 0.43 0.35 / / / 0.62 0.68 0.32
Multi-GraspLLM-split 0.41 1.04 0.32 0.48 0.94 0.41 0.56 0.38 0.31 0.45 0.68 0.38 0.25 0.40 0.45 0.43 0.69 0.37
Multi-GraspLLM-mix 0.37 1.00 0.34 0.42 0.75 0.43 0.48 0.27 0.33 0.34 0.41 0.43 0.20 0.40 0.48 0.36 0.57 0.40
Table 2: Comparison of SOTA methods and Multi-GraspLLM. GroundTruth stands for the preformance of the raw training dataset. “-mix” denotes models trained on mixed datasets, while “-split” indicates models trained separately for each robotic hand. “Avg” stands for the average performance of all hands.
5.1 Dataset and Evaluation Metrics
@@
Multi-GraspLLM is evaluated in two key aspects: grasp intention and physical stability. To evaluate grasp intention accuracy, we employ Chamfer distance (CD) (cm), which quantifies the spatial alignment between the predicted and ground truth robotics hand point clouds. This metric measures how well our inferred grasps match the intended grasp locations on objects, where lower values indicate better alignment. To evaluate grasp quality, we employ three metrics: the maximum penetration distance (Pen) (cm) between the hand and target object measures physical feasibility, the grasp success rate (Suc) in Isaac Sim. More details are in our supplementary material.
@@ 5.2 Implementation Details
For constructing Multi-GraspSet, we employ DexGraspNet [49], GraspIt [28] and Contact-GraspNet [41] as our grasp generation methods. For DexGraspNet [49], we generate 128 grasp poses for the Allegro Hand [53] and 160 for the Shadow Hand [38] per object, using Adam optimizer with 6000 epochs. For GraspIt [28], we generate 30 grasps for Jaco hand [11] and 60 grasps for Barrett hand [2]. For Contact-GraspNet [41], we quickly sample 256 points from each point cloud and generate 200 grasps.
During training, we perform multimodal alignment for 3 epochs with a learning rate of 2e-3, followed by instruction tuning for another 3 epochs with a learning rate of 2e-5. For each object, we uniformly sample 8,192 points from its mesh to generate the point cloud. The grasp angles are discretized into 384 bins, with their upper and lower bounds for different robotic hands determined by their entire dataset.
@@ 5.3 Comparison with State-of-the-Art Methods
We compared our Multi-GraspLLM method with existing open-source SOTA approaches. Since there is currently no open-source method for multi-dexterous hand and gripper grasping generation, we conducted separate comparisons on corresponding end-effectors: SceneDiffuser [9], DexGYS [52] and DexGraspNet [49] were evaluated on dexterous hands, both trained on our dataset, while Contact-GraspNet [41] was evaluated on parallel grippers. The main results are shown in Table 2, “Multi-GraspLLM-split” means that our Multi-GraspLLM model is trained separately on the dataset for each manipulator, while “-mix” refers to mixed training. The results show that mixed training outperforms separate training by a large margin, indicating that joint training can significantly improve model performance. Our methods trained separately or mixed outperform all existing baselines, especially in intention accuracy, with better physical stability and less penetration. Unlike DexGraspNet [49], which adds specific loss functions for physical stability, and the two-stage method DexGYS [52], our approach is end-to-end without any extra loss functions. Our approach outperforms all existing methods and surpasses our original Multi-GraspSet dataset, as demonstrated by the “GroundTruth” presented in Table 2.
Figure 6 shows that Multi-GraspLLM can generate grasp poses based on different contact information: the low-level yields grasps without contact data, the mid-level uses object part information and the high-level leverages finger contact for fine control.
@@ 5.4 Ablation Study
To verify the effectiveness of each component of our method, we performed ablation studies using the average performance of all hands, investigating the impact of each special token and the presence of two-stage training. We also explored the effect of grasp bin length and investigated how the structure of training data affects the model’s performance.
@@
Multi-GraspLLM 0.36 0.57 0.40
w/o hand token 0.38 0.63 0.38
w/o scale token 0.42 0.71 0.35
w/o grasp token 0.41 0.62 0.36
w/o 2-stage training 0.48 0.74 0.32
@@
We first performed ablation on each special token and two-stage training. As shown in Table 3, the two-stage training and the grasp bin token significantly improve performance. A single-stage model struggles to respond effectively to newly introduced grasp bin tokens never encountered before. Due to the incomplete encoding vocabulary for numbers in Vicuna’s [5] tokenizer, larger numbers without grasp bin token such as “400” may be decoded as separate tokens [“40”,“0”], which further hinders the model’s grasp bin comprehension. Additionally, since PointBERT [60] normalizes point clouds and thus missing overall scale information, a scale token helps the model better understand the shape of the point clouds. Lastly, the robotic hand special token still exerts a significant impact, though we only use five types of robotic hands, which the model can memorize by name.
@@
Next, we investigated the effect of the number of grasp bins. The choice of bin number presents a trade-off: larger bin numbers (resulting in smaller bin widths) provide higher resolution, but make prediction more challenging. Smaller bin numbers simplify prediction but reduce the precision of grasp generation. We therefore selected a moderate bin width to balance these competing factors. Moreover, our method with all these bin number configurations consistently outperforms the existing methods (Table 2).
[…]
0.40 0.67 0.37
0.39 0.62 0.35
0.36 0.57 0.40
Table 5: Ablation on training data for instruction tuning.
We then performed an ablation study on the type of data for instruction tuning. As described in Sec. 4.2, our instruction tuning data consists of three types,“Single-Grasp”:“Single-Round Grasp Generation”, “Mix”: “Multi-Round Mix” and “Multi-Grasp”: “Multi-Round Grasp Generation”. As shown in Table 5, fine-tuned on data with more complex and multi-round conversations from “Single-Grasp”,“Mix” to “Multi-Grasp”, our model can achieve consistently better performance. Conceptually, training data with more complex structure allows the model to gain better performance through multitask learning. The introduction of auxiliary tasks (i.e. caption task) prevents the network from exclusively focusing on the primary task, which reduces the risk of overfitting. This training strategy also ensures that the model does not excessively concentrate on the processing of input language while neglecting geometric features of input point clouds, ultimately leading to improved grasp prediction.
6 Conclusion
In this work, we present the first multi-hand grasp dataset with rich semantic guidance. Leveraging this dataset, we propose Multi-GraspLLM, a multimodal LLM method for multi-hand grasp generation. Our method can generate grasp poses for different robotic hands based on linguistic grasp descriptions. Our approach fills a gap in the cross-hand semantic grasp generation field. In the future, we plan to collect larger data using more types of robotic hands and objects, which can further enhance the ability of Multi-GraspLLM to generalize to real-world scenarios, ultimately driving advancements in multi-robot collaboration and human-robot interaction.
Appendix A Additional Implementation Details
A.1 Dataset Annotation
As described in our main paper, we used GPT-4o [1] to generate our basic conversation dataset. Specifically, we developed an LLM-assisted annotation method, which is based on well-crafted instruction prompts to generate natural and diverse conversations. In the following, we provide details on our carefully designed system prompt.
To generate diverse conversations based on different grasp contact information for each object, we adopted a template-based approach. We first used LLM [1] to generate rough templates, then filled these templates with the information of a specific grasp case, and finally used LLM [1] again to refine each dialogue, ensuring diverse and fluent conversations.
- USER: You now need to generate more instruction templates based on a series of templates that I provide to you. These templates focus on asking questions or giving instructions about grasping from the perspective of a robotic hand. Each new template should include “object_name”, “contact_info”, and “robotics_hand”. Below are the templates I provided to you. You need to create more templates similar to these, focusing on different grasping perspectives: templates = [ f“How can the {robotics_hand} securely grasp the {object_name} when {contact_info}?”, f“What is the ideal grasping pose for the {robotics_hand} when {contact_info} with the {object_name}?” ] The output format should be: templates = […]
Table 6: System prompt of generating initial rough template.
- USER: Polish the following sentence to make it smooth and fluent, then output the polished sentence. The general meaning of the sentence should not be altered.“{prompt}” Do not change the word “{robotics_hand}”, and directly output the polished sentence without quotation marks.
Table 7: System prompt of final refinement.
For the initial rough template generation, we started with several manually crafted prompts and used GPT-4 [1] to expand them, as shown in Table 6. These templates were designed to incorporate essential information, such as robotic hand configurations and grasp details, that vary according to different levels of instruction. Each prompt was structured to query the grasp poses or specify the grasp commands for the target objects as shown in Figure 10. We then populated these templates with contact annotations from Multi-GraspSet. Due to the complexity of the contact information involved, we further refined the populated templates using GPT-4 [1] to create the final set of basic conversations, as detailed in Table 7. These iterative refinements helped ensure both clarity and smoothness in the generated dialogue.
A.2 Collection of Multi-GraspSet
We collected 1800 objects from Oakink [58] and 300 objects from ShapeNet [3] as our dataset base. To facilitate reliable intersection testing, we performed convex decomposition using COACD [51] with a threshold of 0.01 to make the meshes watertight. For grasp pose generation, we used DexGraspNet [49], GraspIt [28] and Contact-GraspNet [41] with their default parameter settings. After generating grasp poses, we perform penetration checking on all grasp poses, filtering out invalid candidates where penetration distances exceed the threshold (2cm).
For dexterous hands, each object typically corresponds to numerous possible contact patterns due to the complex structure of the dextrous hand. When constructing the conversation dataset, we randomly selected only one grasp pose for each contact pattern. In contrast, grippers have much simpler mechanics, resulting in only a few possible contact patterns per object. To increase the diversity of gripper grasp pose in our dataset, we randomly selected multiple different grasp poses for each contact information pattern.
Method Low-level mid-level high-level Avg
[…]
DexGraspNet [49] 1.05 0.79 0.37 0.99 0.76 0.29 1.02 0.79 0.34 1.02 0.78 0.33
SceneDiffuser [9] 1.09 0.77 0.31 1.19 0.72 0.36 1.14 0.79 0.26 1.14 0.76 0.31
DexGYS [52] / / / / / / 0.62 0.68 0.32 0.62 0.68 0.32
Contact-GraspNet [41] 0.60 0.41 0.44 0.63 0.46 0.48 / / / 0.62 0.43 0.46
Multi-GraspLLM-mix 0.78 0.64 0.43 0.34 0.51 0.35 0.31 0.57 0.40 0.36 0.57 0.40
Table 8: Comparison of baseline method and Multi-GraspLLM in different instruction levels.
A.3 Multi-GraspLLM
We employ PointBERT [60] as our point cloud encoder, which utilizes the ViT-L/14 model from OpenCLIP [10]. For training data, we structure our 1M basic conversations into two parts: 500k conversations for multimodal alignment based on “Single-Round Grasp Generation”, and another 300k split equally between “Single-Round Grasp Generation”, “Multi-Round Mix”, and “Multi-Round Grasp Generation” for instruction tuning as mentioned in our main paper. We conduct full-parameter fine-tuning of the LLM backbone and modality adaptor using 8 $\times$ 80G A100 GPUs, while testing is performed on a single A100 GPU.
A.4 Evaluation Metrics
As described in our main paper, we evaluate grasp intention and physical stability using multiple metrics: Chamfer distance, maximum penetration distance, grasp success rate. The Chamfer distance is implemented using PyTorch3D [36]. For our calculations, we uniformly sample 512 points on each robotic hand and utilize the point-to-surface distance through kaolin [36] to calculate the maximum penetration distance.
Appendix B Additional Ablation Experiments
We conducted additional experiments on model parameters and instruction levels.
B.1 Effect of Instruction Levels
As described in the main paper, we establish a three-level instruction hierarchy: low-level instructions without contact information, mid-level instructions specifying only the target grasping part, and high-level instructions containing fine-grained grasp information. Specifically, dextrous hands [53, 38] consist of all three instruction levels, while the Panda gripper [7], due to its limited degrees of freedom, only operates at mid and low instruction levels. As shown in Table 8, Multi-GraspLLM showed distinct differences between different instruction levels, whereas our baseline method without semantic guidance achieves consistent performance.
Our method showed varying performance in terms of Chamfer Distance (CD) which represents grasp intention. Multi-GraspLLM achieved the highest CD in low-level instructions that contain no contact information. For mid-level instructions, our model obtained intermediate CD values. Since high-level instructions contain the most detailed contact information, our model demonstrated the lowest CD, which also indicates the best grasp intention consistency. Notably, other grasp quality metrics remained consistent across all levels. These results align with our intuitive expectations.
@@ B.1 Effect of Instruction Levels
B.2 Effect of Training Data Amount for Modality Alignment
We also analyzed the effect of the size of the training data in modality alignment and structure of data in instruction tuning, as shown in Figure 7 and Table 5. We gradually reduced the amount of modality alignment data to test our alignment method. As shown in Figure 7, using more data improves the performance of our model in intention understanding (lower CD) and grasp quality (higher Suc).
@@ B.2 Effect of Training Data Amount for Modality Alignment
One hand 0.41 1.14 0.32
Two hands 0.43 1.23 0.29
Three hands 0.41 1.11 0.31
Four hands 0.39 1.02 0.33
Five hands 0.37 1.00 0.34
Table 9: Performance of Allegro hand in different number of hands used in training.
Figure 8: Robot experiment setup for Allegro.
B.3 Effect of Used Hands in Training
Finally, we analyzed the impact of the used hands involving in training. We evaluated the performance of the Allegro hand as the total number of training hands changed. As shown in Table 9, “Two hands” stands for Allegro and Shadow, “Three hands” means Allegro, Shadow, Panda, “Four hands” means Allegro, Shadow, Panda, and Barrett and “Five hands” means Allegro, Shadow, Panda, Barrett, and Jaco. Starting from three hands, the mixed training results were already better than the separate training results. Moreover, as the number of hands increases and the data volume grows, the model’s performance improves.
Appendix C Real World Experiments
We conducted extensive experiments in real world environments.
C.1 Robot System
As shown in the Figure 8, We built two systems in a real environment to test the model’s performance on a gripper and a multi-fingered dexterous hand. As shown in the left part of Figure 8, our robot gripper setup consists of an AR5X arm with a gripper and an Orbbec DaBai depth camera. We operate the robotic arm at different positions and capture various depth images using the camera at the wrist. Our dexterous hand system is composed of a UR5 arm and an Allegro hand, with two Orbbec Femto Bolt depth cameras, as shown in the right part of Figure 8. In both systems, we obtain full object point clouds from the depth images with the help of SAM2 [37].
We collect 29 objects from daily life covering various objects from tools to potted plants. We tested both the baseline method for grippers [41, 29] and Multi-GraspLLM on these objects, evaluating them based on grasping success rate (Suc) and part selection accuracy (Acc). For each method, we conducted 20 tests and calculated the average performance, with both part selection correctness and grasping success being assessed manually. For both systems, we filtered out all grasp poses that have collisions with the table for evaluation.
method Gripper Allegro
Suc Acc Suc Acc
Contact-GraspNet [41] 0.48 0.42 / /
Scenediffers [9] / / 0.24 0.30
DexGraspNet [49] / / 0.29 0.32
DexGYS [52] / / 0.27 0.56
Multi-GraspLLM 0.45 0.82 0.31 0.72
Table 10: Result of real world experiments.
C.2 Experiments Results
As shown in Table 10, our Multi-GraspLLM achieves comparable grasp success rates to traditional methods that focus on grasp quality and DexGYS [52] that use semantic guidance, while significantly outperforming them in part selection accuracy, which measures grasp intention. These results demonstrate both the sim-to-real feasibility of our approach and its superiority in grasp intention understanding. We show the visualization of grasping experiments under different grasp descriptions in Figure 9.
Figure 9: Illustration of grasp process in real world.
Figure 10: Additional visualization of the grasp pose generated by Multi-GraspLLM.